# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 1.6 Outline of the paper
In Section 4, we will prove all-genus Seiberg duality conjecture for any QP in $\Omega_{n}$ that is mutation-equivalent to $A_{n}$ by utilizing localization Theorem. We will describe the torus fixed points, one-dimensional orbits, and contributions of decorated graphs. Then, we will compare the contribution of each decorated graph to the Gromov-Witten invariants of two quiver varieties in $\Omega_{n}$ that are related by a quiver mutation.
In Section 5, we will prove the cluster algebra conjecture for $A_{n}$-type quivers by using abelian/nonabelian correspondence based on [CFKS08, Web23, Web24].
@@ Acknowledgement
The first author thanks Changzheng Li, Jiayu Song, Yaoxiong Wen, and Mingzhi Yang for many helpful discussions on quantum cohomology rings of flag varieties. Part of the progress was made when the first author was visiting BIMSA, YMSC, and IASM; he is grateful for their hospitality. The second author expresses sincere gratitude to Peng Zhao for suggesting a possible connection between Gromov-Witten theory and cluster algebras. She also thanks Professor Yongbin Ruan for his valuable discussions, encouragement to complete this work, and for providing a supportive and enjoyable working environment. She is also thankful to Wei Gu for explaining his work about nonabelian mirror symmetry. Special thanks go to Professor Fang Li for helpful discussions on cluster algebras, and to Professor Sergey Fomin, Jiarui Fei, Fan Qin and Zijun Zhou for their interest in this work and for suggesting related problems. Most of the work was carried out while the second author was at Zhejiang University, and she is deeply grateful for the support provided by the university. The first author is supported by the National Key R&D Program of China (Project No. 2023YFA1009802) and by NSFC. 12422104. The second author is supported by NSFC.12301080.
@@ 2.1 Quivers with potentials
  • Replace the factor $ab$ in the potential function $W$ by the new one $c$ and add all cubic terms $cb^{*}a^{*}$ to the potential. Denote the resulting potential function by $W^{\prime}$, and denote the resulting QP by $\tilde{\mathbf{Q}}=(Q_{f}\subset Q_{0},\tilde{Q}_{1},\tilde{W})$.
  • Let ${\mathbf{Q}^{\prime}}=(Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime})$ be the reduced QP of $\tilde{\mathbf{Q}}$, see [DWZ08, Theorem 4.6]. This is the new QP we get by performing the quiver mutation $\mu_{v}$ to the original $\mathbf{Q}$.
We call two QPs mutation-equivalent if we can obtain one by performing a sequence of quiver mutations to the other one. For all examples in this work, $\mathbf{Q}^{\prime}=\tilde{\mathbf{Q}}$. Note that the set of nodes are preserved by quiver mutations up to order.
@@
Figure 1: The dashed arrows in the left QP are canceled. The right one has a potential function $W=bac$.
@@ 2.2 Cluster algebras
We introduce the definition of cluster algebras of geometric type. We refer readers to the original works for the theory of cluster algebras [FZ02, FZ03, BFZ05, FZ07, DWZ08, DWZ10] and the nice introductory book [FWZ21c, FWZ21d, FWZ21a, FWZ21b] which can be found in the portal of Fomin.
@@
  • $\mathbf{x}=(x_{1},\ldots,x_{n})$ is a set of n algebraically independently variables of $\mathscr{F}$ such that $\mathscr{F}=\mathbb{Q}(x_{1},\ldots,x_{n})$.
The variable set $\mathbf{x}$ is called the cluster and variables $x_{1},\ldots,x_{n}$ are called cluster variables of the seed $(\mathbf{Q},\mathbf{x})$.
@@
    $x_{v}x_{v}^{\prime}=\prod_{u\in Q_{0}}x_{u}^{[b_{uv}]_{+}}+\prod_{w\in Q_{0}}x_{w}^{[b_{vw}]_{+}}\,.$ (2.1)
@@
Notice that we work with unlabeled seeds, which means we identify two seeds $(\mathbf{Q},\mathbf{x})$ and $(\mathbf{Q}^{\prime},\mathbf{x}^{\prime})$ if the QPs $\mathbf{Q}$ and $\mathbf{Q}^{\prime}$ differ by relabeling their nodes and clusters $\mathbf{x}$ and $\mathbf{x}^{\prime}$ differ by a permutation.
@@
Figure 2:
@@
$(\frac{x_{1}+x_{3}}{x_{2}},\frac{x_{1}+x_{3}+x_{2}x_{3}}{x_{1}x_{2}},x_{3})\xrightarrow[]{\mu_{2}}(\frac{x_{1}+x_{3}}{x_{2}},x_{1},x_{3})\xrightarrow[]{\mu_{1}}(x_{2},x_{1},x_{3}).$ (2.2)
@@
In the above example, $(1)$ all non-initial cluster variables can be written as rational functions of the initial variables whose denominators are monomials of initial cluster variables and $(2)$ numerators are polynomials of initial cluster variables with positive integer coefficients. The two observation $(1)$ and $(2)$ are called the Laurent phenomenon and the positivity conjecture in cluster algebra theory. The Laurent phenomenon is proved in [FZ02] where the positivity conjecture is proposed. The positivity conjecture is proved in [KQ14, MSW11, LS15].
2.3 $A_{n}$-mutation-equivalent quivers
@@
The work [BV08] has given a precise description of quivers that are mutation-equivalent to the $A_{n}$ quiver without frozen nodes. Our situation is similar with that in [BV08] with slight modifications. Let $\Omega_{n}^{A}$ denote the set of QPs that are mutation-equivalent to the $A_{n}$-type quiver in Figure 3. Let $vf$ denote the only frozen node, which we refer to as the root of the quiver.
Lemma 2.9 ([BV08]).
@@
  • A gauge node has at most four adjacent arrows. If a gauge node $v$ has only two arrows, then it can be a sink, source or there is a path passing through $v$. If a gauge has exactly three adjacent arrows, then two of them are contained in a 3-cycle, and the third one doesn’t belong to any 3-cycle. If a gauge node has four adjacent arrows, then two of them belong to a 3-cycle and the remaining two belong to another 3-cycle.
@@ 2.4 Quiver varieties
Now we move to the definition of quiver varieties. We refer readers to the nice book [Kir16]. Fix a QP $\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W)$, and decorate it by integers $\mathbf{r}=(r_{v})_{v\in Q_{0}}$.
@@
    $(g_{v})_{v\in Q_{0}\backslash Q_{f}}\cdot(A_{s(a)\rightarrow t(a)})_{a\in Q_{1}}=(g_{s(a)}A_{s(a)\rightarrow t(a)}g_{t(a)}^{-1})_{a\in Q_{1}},$ (2.3)
@@
In the above definition we always assume that $V^{ss}_{\theta}(G)=V^{s}_{\theta}(G)$ and $Z^{ss}_{\theta}(G)=Z^{s}_{\theta}(G)$.
@@
For a decorated QP $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, define the outgoing of a gauge node $v$ to be $N_{f}(v)=\sum_{a\in Q_{1},s(a)=v}r_{t(a)}$, and the incoming of the node $v$ to be $N_{a}(v)=\sum_{a\in Q_{1},t(a)=v}r_{s(a)}$.
@@
The input data for the QP includes $V=\oplus_{i=1}^{n}\mathbb{C}^{N_{i}\times N_{i+1}}$, $G=\prod_{i=1}^{n}GL(N_{i})$, and $G$ acts on $V$ as (2.3). Choose the character $\theta=\prod_{i=1}^{n}\det(g_{i})^{\sigma_{i}}$ for $\sigma_{i}>0$. Then
@@ 2.5 Quiver mutations for decorated quivers
$r_{u}^{\prime}=\begin{cases}\max\{N_{f}(v),N_{a}(v)\}-r_{v}&\text{ if }u=v\\ r_{u}&\text{ otherwise}.\end{cases}$ (2.4)
@@
  • Step one, we start from the root $v_{f}$ whose assigned integer is always $N_{n+1}$. By Lemma 2.10, there are at most $2$ nodes attached to it which are denoted by $v_{1},v_{2}$ with arrows being shown in Figure 4. The integer assigned to $v_{1}$ is $r_{v_{1}}=N_{l}$ and the integer assigned to $v_{2}$ is $r_{v_{2}}=N_{n+1}-N_{l+1}$ for $n\geq l\geq 0$. We formally set $N_{0}=0$. When $l=0$, then there is only one node $v_{2}$ attached to $v_{f}$, and when $l=n$, then $N_{n+1}-N_{l}=0$ and there is only one node $v_{1}$ attached to $v_{f}$, see Figure 4 $(2)$ and $(3)$.
  • Step two, if there are exactly two nodes $v_{1}\rightarrow v_{f}\rightarrow v_{2}\rightarrow v_{1}$, $\mathbf{Q}\backslash\{v_{f},v_{2}\rightarrow v_{1}\}=\mathbf{Q}_{1}\sqcup\mathbf{Q}_{2}$. The subquiver $\mathbf{Q}_{1}$ lies in $\Omega^{A}_{l-1}$ and is mutation-equivalent to the $A_{l-1}$-type quiver with decorations $N_{1}<N_{2}<\cdots<N_{l}$, and $\mathbf{Q}_{2}$ is mutation-equivalent to $A_{n-l-1}$ with decorations $N_{l+2}-N_{l+1}<\cdots<N_{n+1}-N_{l+1}$. The cases when there is only one node $v_{1}\,(v_{2})$adjacent to the $v_{f}$ can be obtained by letting $l=n\,(l=0)$.
  • Step three, recursively construct the decorations of the quivers $\mathbf{Q}_{i}$ in Step two by utilizing the step one and two.
@@
From the above recursive relation, one can find that if $v\prec v^{\prime}$, the $r_{v}<r_{v^{\prime}}$. One can also easily get the for any node $v\in Q_{0}$, the decorated integer can only be $r_{v}=N_{k}-N_{l}$ for some $0\leq l<k$.
By abuse of notations, we still denote the set of decorated QPs that are mutation equivalent to the decorated $A_{n}$ by $\Omega^{A}_{n}$.
@@
  • When the node $v$ has exactly four adjacent arrows, which we denote by $v_{1},\cdots,v_{4}$, the local picture can be described as Figure 6 (a) for $m>k>p>l\geq 0$. We formally let $N_{0}=0$.
    The case when $N_{f}(v)<N_{a}(v)$ is on the right-hand side and one can find it is mutation equivalent to the left-hand side one.
  • The case when the node $v$ has only three adjacent nodes can be realized by taking special values for $m>k>p>l\geq 0$.
    • When $p=l$, $N_{p}-N_{l}=0$. In this case, we delete the node $v_{3}$, and the node $v$ has two arrows going out and one arrow coming in.
    • When $k=m-1$, $N_{m}-N_{k+1}=0$. We then delete the node $v_{2}$ and the node $v$ will only have three adjacent arrows.
@@
  • Similarly, the case when the node $v$ has only two adjacent arrows can be obtained by taking special values for $m,k,p,l$. When $p=l,k=l+1$, the node $v$ has two adjacent arrows which form a three-cycle. When $m=k+1,p=l$, the node $v$ is a source and has two adjacent arrows going out.
  • When $m=k+1=l+2$, the node $v$ has only one adjacent arrow.
@@
2.6 Varieties of QPs in $\Omega^{A}_{n}$
Let $(\mathbf{Q}=(Q_{0},Q_{1},W),\mathbf{r})\in\Omega^{A}_{n}$ be an arbitrary decorated QP in $\Omega^{A}_{n}$. Denote the input data of the quiver by $(V,G,\theta)$ where $G=\prod_{i\in Q_{0}\backslash Q_{f}}GL(r_{i})$ and $\theta(g)=\prod_{i\in Q_{0}\backslash Q_{f}}\det(g_{i})^{\sigma_{i}}$.

### v2

@@ 1.6 Outline of the paper
In Section 4, we will prove the all-genus Seiberg duality conjecture for any QP in $\Omega_{n}$ that is mutation-equivalent to $A_{n}$ by utilizing the localization theorem. We will describe the torus fixed points, one-dimensional orbits, and contributions of decorated graphs. Then, we will compare the contribution of each decorated graph to the Gromov-Witten invariants of two quiver varieties in $\Omega_{n}$ that are related by a quiver mutation.
In Section 5, we will prove the cluster algebra conjecture for $A_{n}$-type quivers by using abelian/nonabelian correspondence based on [abelian/nonabelian:CKB, abeliannonabelian:Webb, abelianizationlef:Webb].
@@ Acknowledgement
The first author thanks Changzheng Li, Jiayu Song, Yaoxiong Wen, and Mingzhi Yang for many helpful discussions on quantum cohomology rings of flag varieties. Part of the progress was made when the first author was visiting BIMSA, YMSC, and IASM; he is grateful for their hospitality. The second author expresses sincere gratitude to Peng Zhao for suggesting a possible connection between Gromov-Witten theory and cluster algebras. She also thanks Professor Yongbin Ruan for his valuable discussions, encouragement to complete this work, and for providing a supportive and enjoyable working environment. She is also thankful to Wei Gu, Yaoxiong Wen, Jie Zhou, and Zijun Zhou for discussions on nonabelian mirror symmetry and quantum cohomology of nonabelian GIT quotients. Special thanks go to Professor Fang Li for helpful discussions on cluster algebras, and to Professor Sergey Fomin, Jiarui Fei, Fan Qin, and Cheng Shu for their interest in this work and for suggesting related problems. Most of the work was carried out while the second author was at Zhejiang University, and she is deeply grateful for the support provided by the university. The first author is supported by the National Key R&D Program of China (Project No. 2023YFA1009802) and by NSFC. 12422104. The second author is supported by NSFC.12301080.
@@ 2.1 Quivers with potentials
  3. 3.
    Replace the factor $ab$ in the potential function $W$ by the new one $c$ and add all cubic terms $cb^{*}a^{*}$ to the potential. Denote the resulting potential function by $\tilde{W}$, and denote the resulting QP by $\tilde{\mathbf{Q}}=(Q_{f}\subset Q_{0},\tilde{Q}_{1},\tilde{W})$.
  4. 4.
    Let ${\mathbf{Q}^{\prime}}=(Q_{f}\subset Q_{0},Q_{1}^{\prime},W^{\prime})$ be the reduced QP of $\tilde{\mathbf{Q}}$, see [clusterpot1, Theorem 4.6]. This is the new QP we get by performing the quiver mutation $\mu_{v}$ to the original $\mathbf{Q}$.
We call two QPs mutation-equivalent if we can obtain one by performing a sequence of quiver mutations to the other one. For all examples in this work that are mutation equivalent to the $A_{n}$-type quiver, $\mathbf{Q}^{\prime}=\tilde{\mathbf{Q}}$. Note that the set of nodes is preserved by quiver mutations up to order.
@@
Figure 1: The two quiver with potentials are related by a quiver mutation at the center node. The dashed arrows in the left QP are canceled. The right one has a potential function $W=bac$.
@@ 2.2 Cluster algebras
We introduce the definition of cluster algebras of geometric type. We refer readers to the original works for the theory of cluster algebras [clusteralg:1, clusteralg:2, clusteralg:3, clusteralg:4, clusterpot1, clusterpot:2] and the introductory book [fomin:cluster1-3, fomin:cluster4-5, fomin:cluster6, fomin:cluster7].
@@
  2. 2.
    $\mathbf{x}=(x_{1},\ldots,x_{n})$ is a set of n algebraically independently variables and let $\mathscr{F}=\mathbb{Q}(x_{1},\ldots,x_{n})$.
The set $\mathbf{x}$ is called a cluster and variables $x_{1},\ldots,x_{n}$ are called cluster variables of the seed $(\mathbf{Q},\mathbf{x})$.
@@
    (2) $x_{v}x_{v}^{\prime}=\prod_{u\in Q_{0}}x_{u}^{[b_{uv}]_{+}}+\prod_{w\in Q_{0}}x_{w}^{[b_{vw}]_{+}}\,.$
@@
Notice that we work with unlabeled seeds, which means we identify two seeds $(\mathbf{Q},\mathbf{x})$ and $(\mathbf{Q}^{\prime},\mathbf{x}^{\prime})$ if the QPs and clusters differ by relabeling.
@@
Figure 2: An $A_{2}$ quiver with one frozen node.
@@
$(\frac{x_{1}+x_{3}}{x_{2}},\frac{x_{1}+x_{3}+x_{2}x_{3}}{x_{1}x_{2}},x_{3})\xrightarrow[]{\mu_{2}}(\frac{x_{1}+x_{3}}{x_{2}},x_{1},x_{3})\xrightarrow[]{\mu_{1}}(x_{2},x_{1},x_{3}).$
@@
In the above example, $(1)$ all non-initial cluster variables can be written as rational functions of the initial variables whose denominators are monomials of initial cluster variables and $(2)$ numerators are polynomials of initial cluster variables with positive integer coefficients. The two observations $(1)$ and $(2)$ are called the Laurent phenomenon and the positivity conjecture in cluster algebra theory. The Laurent phenomenon is proved in [clusteralg:1] where the positivity conjecture is proposed. The positivity conjecture is proved in [cluster:positivityQin, cluster:positivityMSL1, cluster:positivityLSR2].
2.3 A-mutation-equivalent quivers
@@
The work [Amutationequivalence] has given a precise description of quivers that are mutation-equivalent to the $A_{n}$ quiver without frozen nodes. Our situation is similar to that in [Amutationequivalence] with some modifications. Let $\Omega_{n}^{A}$ denote the set of QPs that are mutation-equivalent to the $A_{n}$-type quiver in Figure 3. Let $v_{f}$ denote the only frozen node, which we refer to as the root of the quiver.
Lemma 2.9 ([Amutationequivalence]).
@@
  3. 3.
    A gauge node has at most four adjacent arrows. If a gauge node $v$ has only two arrows, then it can be a sink, source or there is a path passing through $v$. If a gauge node has exactly three adjacent arrows, then two of them are contained in a 3-cycle, and the third one doesn’t belong to any 3-cycle. If a gauge node has four adjacent arrows, then two of them belong to a 3-cycle and the remaining two belong to another 3-cycle.
  4. 4.
@@ 2.4 Quiver varieties
Now we move to the definition of quiver varieties. We refer readers to the book [quiver]. Fix a QP $\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W)$, and decorate it with integers $\mathbf{r}=(r_{v})_{v\in Q_{0}}$.
@@
    (3) $(g_{v})_{v\in Q_{0}\backslash Q_{f}}\cdot(A_{s(a)\rightarrow t(a)})_{a\in Q_{1}}=(g_{s(a)}A_{s(a)\rightarrow t(a)}g_{t(a)}^{-1})_{a\in Q_{1}},$
@@
In the above definition, we always assume that $V^{ss}_{\theta}(G)=V^{s}_{\theta}(G)$ and $Z^{ss}_{\theta}(G)=Z^{s}_{\theta}(G)$.
@@
For a decorated QP $(\mathbf{Q}=(Q_{f}\subset Q_{0},Q_{1},W),\mathbf{r})$, define the outgoing of a gauge node $v$ to be $N_{f}(v)=\sum_{a\in Q_{1},s(a)=v}r_{t(a)}$, and the incoming to be $N_{a}(v)=\sum_{a\in Q_{1},t(a)=v}r_{s(a)}$.
@@
The input data for the QP includes $V=\oplus_{i=1}^{n}\mathbb{C}^{N_{i}\times N_{i+1}}$, $G=\prod_{i=1}^{n}GL(N_{i})$, and $G$ acts on $V$ as (3). Choose the character $\theta=\prod_{i=1}^{n}\det(g_{i})^{\sigma_{i}}$ for $\sigma_{i}>0$. Then
@@ 2.5 Quiver mutations for decorated quivers
(4) $r_{u}^{\prime}=\begin{cases}\max\{N_{f}(v),N_{a}(v)\}-r_{v}&\text{ if }u=v\\ r_{u}&\text{ otherwise}.\end{cases}$
@@
  1. 1.
    Step one, we start from the root $v_{f}$ whose assigned integer is always $N_{n+1}$. By Lemma 2.10, there are at most $2$ nodes attached to it which are denoted by $v_{1},v_{2}$ with arrows being shown in Figure 4. The integer assigned to $v_{1}$ is $r_{v_{1}}=N_{l}$ and the integer assigned to $v_{2}$ is $r_{v_{2}}=N_{n+1}-N_{l+1}$ for $n\geq l\geq 0$. We formally set $N_{0}=0$. When $l=0$, then there is only one node $v_{2}$ attached to $v_{f}$. When $l=n$, then $N_{n+1}-N_{l}=0$ and there is only one node $v_{1}$ attached to $v_{f}$, see Figure 4 $(2)$ and $(3)$.
  2. 2.
    Step two, if there are exactly two nodes $v_{1}\rightarrow v_{f}\rightarrow v_{2}\rightarrow v_{1}$, $\mathbf{Q}\backslash\{v_{f},v_{2}\rightarrow v_{1}\}=\mathbf{Q}_{1}\sqcup\mathbf{Q}_{2}$. The subquiver $\mathbf{Q}_{1}$ lies in $\Omega^{A}_{l-1}$ and is mutation-equivalent to the $A_{l-1}$-type quiver with decorations $N_{1}<N_{2}<\cdots<N_{l}$. The other one $\mathbf{Q}_{2}$ is mutation-equivalent to $A_{n-l-1}$ with decorations $N_{l+2}-N_{l+1}<\cdots<N_{n+1}-N_{l+1}$. The cases when there is only one node $v_{1}\,(v_{2})$ adjacent to the $v_{f}$ can be obtained by letting $l=n\,(l=0)$.
  3. 3.
    Step three, recursively construct the decorations of the quivers $\mathbf{Q}_{i}$ in Step two by viewing $v_{i}$ as root of the new quivers $\mathbf{Q}_{i}$ and utilizing the steps one and two.
@@
From the above recursive relation, one can find that if $v\prec v^{\prime}$, the $r_{v}<r_{v^{\prime}}$. One can also easily get that for any node $v\in Q_{0}$, the assigned integer can only be $r_{v}=N_{k}-N_{l}$ for some $10\leq l<k\leq n+$ 1.
By abuse of notation, we still denote the set of decorated QPs that are mutation-equivalent to the decorated $A_{n}$ by $\Omega^{A}_{n}$.
@@
  1. 1.
    When the node $v$ has exactly four adjacent arrows, which we denote by $v_{1},\cdots,v_{4}$, the local picture can be described as Figure 6 (a) for $m-1>k>p+1\geq l\geq 0$. We formally let $N_{0}=0$.
    The case when $N_{f}(v)<N_{a}(v)$ is in Figure 6 (b) and one can find it is mutation-equivalent to the left-hand side one.
  2. 2.
    The case when the node $v$ has only three adjacent nodes can be realized by taking special values for $m,k,p,l$.
    • When $p=l$, $m-1>k>l+1$, then $N_{p}-N_{l}=0$. In this case, we delete the node $v_{3}$, and the node $v$ has $3$ adjacent nodes $v_{1},v_{2},v_{4}$.
@@
    • When $k=m-1$, $N_{m}-N_{k+1}=0$ and the node $v_{2}$ vanishes.
  3. 3.
    The case when the node $v$ has only two adjacent arrows can also be obtained by taking special values for $m,k,p,l$. When $p=l,k=l+1$, the node $v$ has two adjacent nodes $v_{1},v_{2}$ and they form a three-cycle. When $m=k+1,p=k-1$, the node $v$ has two adjacent nodes $v_{1},v_{3}$. When $m=k+1,p=l$, the node $v$ is a source and has two adjacent $v_{1},v_{4}$.
  4. 4.
    When $m=k+1=l+2$, the node $v$ has only one adjacent node.
@@
2.6 Varieties of QPs mutation equivalent to A-type
Let $(\mathbf{Q}=(Q_{0},Q_{1},W),\mathbf{r})\in\Omega^{A}_{n}$ be an arbitrary decorated QP in $\Omega^{A}_{n}$. Denote the corresponding input data by $(V,G,\theta)$ where $G=\prod_{i\in Q_{0}\backslash Q_{f}}GL(r_{i})$ and $\theta(g)=\prod_{i\in Q_{0}\backslash Q_{f}}\det(g_{i})^{\sigma_{i}}$.